The reversible chemical reaction

A+B⇌C+D

has the following equilibrium constant:

Kc=[C][D][A][B]=2.0

Part A

Initially, only A and B are present, each at 2.00 M. What is the final concentration of A once equilibrium is reached?

Express the molar concentration numerically using two significant figures.

Part B

What is the final concentration of D at equilibrium if the initial concentrations are [A] = 1.00 M and [B] = 2.00 M ?

Express the molar concentration numerically using two significant figures.
Answer #1

Af B 2-0 2-0 0 0 (Ini Haly) Cerarge) -XX Ceguidibium) Ce). L0) K (A) (b) 2-0=()n) (2.0-)F2.0-k) M 1172 A= (2-0-) う (2-0-112)

If you have any query please comment

If you satisfied with the solution please like it thankxx